A Ferris wheel with a radius of 9.0 m rotates at a constant rate, completing one revolution every 36 s .
Suppose the Ferris wheel begins to decelerate at the rate of 0.20 rad/s2 when the passenger is at the top of the wheel.

a)Find the magnitude of the passenger's acceleration at that time.

i got 1.8 and that was correct.

b) Find the direction of the passenger's acceleration at that time.
Answer #1

angular acceleration

ar = r (2 pi / T) ^2 = 9* ( 2* 3.14 / 36)^2

ar = 0.2739 m/s^2

tangential acceleration

at = 9 * 0.2 = 1.8 m/s^2

direction

x = arctan ( ar/at) = 8.652 deg

======

Comment before rate in case any doubt, will reply for sure.. goodluck